OpenStudy (anonymous):

Evaluate the integral ∫1->2∫2->3 of (3x+y)^(−2) dydx

OpenStudy (anonymous):

double integral?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\int\limits_{2}^{3}(3x+y)dydx?\]

OpenStudy (anonymous):

its that except the quantity on parens is ^(-2)

OpenStudy (anonymous):

\[\int\limits\limits_{1}^{2}\int\limits\limits_{2}^{3}(3x+y)^{-2} dydx\]

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

would du = dy?

OpenStudy (anonymous):

i have integral from 1 to 2 of [-1/u] evaluated from 2 to 3 dx

OpenStudy (anonymous):